OpenStudy (ashleneperalez):

Which equation has (2,-1) as a solution? Y=2x-1 Y=x+3 Y=x-3 y=-2x+1

OpenStudy (algorithm):

2 and -1 are the roots of the equation right?

OpenStudy (ashleneperalez):

Yes

OpenStudy (luckyness):

Okay so its pretty difficult to explain, but (x,y) < That is how it is. So if its (2,-1) X=2 Y=-1 And now replace with x as well. Now we have to see which one of these equations = -1 So we have to replace and solve: -1=2*2-1=3 ~False -1=2+3 = 5 ~False -1=2-3 = -1 ~True -1=-2*2+1=5 ~False Answer: -1=2-3 Therefore: C